OpenStudy (anonymous):

Line AB contains points A (0, 1) and B (1, 5). The slope of line AB is −4 negative 1 over 4 1 over 4 4

OpenStudy (welshfella):

for any points (x1, y1) and (x2,y2) the slope of the line joining them is (y2-y1) / (x2 - x1)

OpenStudy (anonymous):

which do you think it is ?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

no.. not -4

OpenStudy (welshfella):

let y2 = 5 , y1 = 1, x2 = 1 and x1 = 0 in the above formula and you'll have it

OpenStudy (welshfella):

slope = (5-1)/ (1-0) = ?

OpenStudy (anonymous):

the slope is 4 because you would rise 4 and run 1 if you understand that

OpenStudy (welshfella):

yes rise = 5-1 , run = 1-0

OpenStudy (welshfella):